# Fine-Tuning Masked Diffusion for Provable Self-Correction

### v3

@@ 4.3 Appyling PRISM to 88B MDM
Setup. We take LLaDA-88B-Instruct (nie2025large), an open-sourced 88B MDM. While freezing the backbone, we add an additional head and a LoRA adapter (hu2022lora), for a total of $\approx 250$M trainable parameters. We fine-tune it using the PRISM loss on the opc-sft-stage2 (Huang2024OpenCoderTO) (0.10.1M pairs) for $100$ epochs. Fine-tuning completes in $\approx 30$ hours on $12$ H100 GPUs.
@@
Results. We evaluate the resulting model in a zero-shot setting on MBPP (austin2021program) and HumanEval (chen2021evaluating), a challenging Python coding benchmarks. As shown in Table 2, PRISM outperforms ReMDM and ReMDM-conf across different numbers of sampling steps. We detail training and inference configurations in Appendix D.6D.7.
Sampling steps 256 512 1024
PRISM 28.0% 39.0% 42.7%
ReMDM 26.2% 34.8% 42.5%
ReMDM-conf 25.6% 34.1% 36.6%
MDM 25.6% 34.1% 36.6%
Table 1: HumanEval zero-shot performance.
Sampling steps 256 512 1024
PRISM 21.8% 29.1% 32.3%
ReMDM 19.7% 28.5% 32.9%
ReMDM-conf 20.1% 29.0% 32.5%
MDM 18.2% 27.8% 31.9%
Table 2: MBPP zero-shot performance.
@@ 5 Conclusion
Discussion on experimental results. MDMs model per-position unmasking posteriors rather than the joint distribution over positions. Consequently, with few-step inference, they are prone to dependency errors across positions (liu2024discrete; xu2024energy; kang2025parallelbench), which creates a natural opportunity for our self-correction mechanism to detect erroneous predictions. We confirm this empirically: methods that incorporate remasking outperform vanilla MDM inference as the number of sampling steps decreases (Figure 4, Table 2). In particular, PRISM achieves larger gains than baselines by learning per-token quality. However, as the number of sampling steps increases, partially unmasked sequences contain fewer errors, leaving less headroom for the self-correction. For the scaling-up experiments in Section 4.3, although we fine-tune LLaDA-8B-Instruct on only 0.1M dataset pairs, we expect that a more generalized model could be obtained by constructing a richer and more diverse dataset, given sufficient compute.
@@
A unique feature of the PRISM loss is that it provably learns per-token quality without fully generating a clean sequence. This sheds light on PRISM’s potential compared to other approaches, such as reinforcement learning, where full sequence generation is necessary to obtain a learning signal. That said, our notion of per-token quality score has limits; as it is based on the per-position posterior marginals, it cannot fully capture global errors, e.g., reasoning. An important direction for future work is to develop frameworks that enable discrete diffusion models to detect global reasoning errors.
@@ Appendix A Related Works
Masked Diffusion Models. Masked Diffusion Models (MDMs) were previously built on continuous-time Markov chains over discrete spaces (hoogeboom2021argmax). Subsequently, austin2021structured introduced D3PM with several families of transition kernels, followed by SEDD (lou2023discrete). These discrete diffusions can also be understood through the lens of discrete flows (campbell2024generative; gat2024discrete), which parallel flow matching and stochastic interpolants over continuous spaces (albergo2022building; albergo2023stochastic; lipman2022flow). Among transition-kernel designs, the absorbing state (masking) kernel has been particularly popular due to its simple formulation and training objective. Subsequent work established a theoretical framework for these models (zheng2023reparameterized; sahoo2024simple; zheng2024masked). Follow-up works have scaled MDMs up to billions of parameters, showcasing their potential compared to autoregressive models (nie2025large; dream2025; gemini2025diffusion; labs2025mercury; gong2025diffucoder).
Inference-time strategies for discrete diffusion models. A key feature of MDMs is their inference-time flexibility, highlighted by zheng2024masked; ou2024your and aligned with our any-order perspective in Section 2.1. Recall the MDM inference procedure from Section 2.1:
@@
The choice of $\mathcal{S}$ is entirely flexible, and an appropriate selection can substantially improve downstream task performance. Since our work focuses on self-correction, i.e., re-masking, we only briefly survey prior work on unmasking strategies (kim2025train; peng2025pathplanningmaskeddiffusion; chang2022maskgit; ben2025accelerated; rout2025anchored). A complementary line of work studies inference-time planning strategies for discrete diffusion with uniform transition kernels (liu2024think; varma2024glauber); these methods fall outside the scope of MDMs.
A.1 Details on prior work on remasking strategy (self-correction) for MDM
@@
  • DFM (gat2024discrete) and ReMDM (wang2025remasking) randomly remask clean tokens in $\mathbf{y}$}, which is equivalent to assigning i.i.d. per-token qualities $g_{\star}^{i}({\mathbf{y}})\sim\operatorname{Unif}[0,1]$.
  • P2-Self (peng2025pathplanningmaskeddiffusion) defines per-token quality using the pretrained MDM’s score for the observed token under input $\mathbf{y}$}, e.g., the probability $f_{\theta}^{i}({\mathbf{y}}^{i}\mid{\mathbf{y}})$. Given that ${\mathbf{y}}^{i}\neq\mathbf{m}$, the model was never trained on this index in this state; it only encodes a meaningful quantity (unmasking posterior) when ${\mathbf{y}}^{i}=\mathbf{m}$. This notion of per-token quality is therefore not grounded.
  • ReMDM-conf (wang2025remasking) takes the per-token quality to be the likelihood at the step the token was unmasked: if position $i$ was unmasked at $t_{k}>t_{\ell}$ (i.e., an earlier step), set $g_{\star}^{i}:=f_{\theta}^{i}({\mathbf{x}}_{t_{k}}^{i}\mid{\mathbf{x}}_{t_{k}})=f_{\theta}^{i}({\mathbf{x}}_{t_{\ell}}^{i}\mid{\mathbf{x}}_{t_{k}})$. As noted, this ignores the current state ${\mathbf{x}}_{t_{\ell}}$ and can therefore be misaligned with the model’s present beliefs.
  • Token Critic (lezama2022improved; lezama2022discrete) and P2-train (peng2025pathplanningmaskeddiffusion) train an auxiliary scorer for the likelihood of a given prediction of clean sequence $\hat{{\mathbf{x}}}$. Given ${\mathbf{x}}_{t_{\ell}}$, they first form a full completion $\hat{{\mathbf{x}}}$ by unmasking each mask token at position $i$ from $f_{\theta}^{i}(\cdot\,\mid\,{\mathbf{x}}_{t_{\ell}})$, then evaluate quality by passing this $\hat{{\mathbf{x}}}$ to the external module. A key limitation is that $\hat{{\mathbf{x}}}$ is drawn from per-token unmasking posteriors, not the ground-truth posterior across positions; thus, even with a perfect pretrained MDM, $\hat{{\mathbf{s}}}$ can deviate substantially from the true likelihood ${\mathbf{x}}\sim p_{\mathrm{data}}$, losing a theoretical guarantee.
  • GIDD (von2025generalized) introduces a family of transition kernels that linearly combine uniform and absorbing (masking) kernels. Models pretrained under this kernel encode posteriors at clean-token positions–however, this requires training under the new kernel and is not directly applicable to an off-the-shelf pretrained MDM, altogether losing the benefit of MDM over other types of kernels.
  • Informed corrector (zhao2024informed) employs a hollow-transformer $g_{\phi^{\prime}}$ satisfying $g_{\phi^{\prime}}^{i}({\mathbf{y}})=g_{\phi^{\prime}}^{i}({\mathbf{y}}\oplus\mathbf{m}_{i})\approx p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\mid{\mathbf{y}}\oplus\mathbf{m}_{i})$, enabling the same training loss as MDM. This, however, mandates a modified architecture (attention tweak) that processes $\mathbf{y}$} as if position $i$ were masked, necessitating retraining.
@@ B.2 Marginal Equivalence of the Masking Process
We formalize that the two masking procedures stated in Section 2.1 induce the same conditional distribution of $\mathbf{z}$} given $\mathbf{x}$}, hence the same joint distributions $({\mathbf{x}},{\mathbf{z}})$. It suffices to show they generate the same distribution over masked index sets.
Consider (b): draw $t\sim\mathrm{Unif}[0,1]$ and for each $i$, replace ${\mathbf{x}}^{i}\in\mathcal{V}$ with $\mathbf{m}$ with probability $t$. The probability of $n$ tokens to be masked is:
@@
where $\mathrm{B}$ denotes Beta function. This coincides with procedure (a), where $n$ is uniformly sampled from $\{0,1,\dots,L\}$.
@@ B.3 PRISM provably learns the per-token quality
In this section, we restate Proposition 1 and detail its proof. The proof essentially follows Section 3.1, but we provide the detailed calculations here.
Proposition 1 (restated). Let the PRISM loss be
@@ B.4 Extension on PRISM’s provable guarantee
In this section, we provide the pseudocode of the practically used PRISM fine-tuning pipeline and then state its theoretical guarantee.
We note that the choice of index set $\mathcal{S}$ at line 5 of Algorithm 1 is flexible; it can either be random or confidence-based. To clarify, at line 6, although we use $f_{\theta}$ to obtain $\mathbf{y}$}, we use the stop-gradient operation.
@@
Proposition 2 (Provable guarantee of PRISM-extension).
@@
By using this property, we now complete the proof; the remaining parts are the same as Proposition 1.
@@
Discussion on a new notion of per-token quality. Proposition 2 guarantees that the minimizer of the extended PRISM loss satisfies
@@
Note that by construction, we always have $i\in\mathcal{S}$. This encompasses the $k=1$ case in Section 3.1; When $k=1$, the posterior distribution $\Pi(\cdot\,|\,{\mathbf{y}},i)$ collapses to a point mass at $\mathcal{S}=\{i\}$, recovering Proposition 1.
However, we emphasize that this extended definition still captures our intended notion of per-token quality—namely, the (posterior) likelihood of ${\mathbf{y}}^{i}$ given the rest of the context, here defined under a certain distribution over ${\mathbf{y}}\oplus\mathbf{m}_{\mathcal{S}}$. (note that always $i\in\mathcal{S}$). Consequently, a model fine-tuned with the (extended) PRISM loss retains the capacity to identify low-quality tokens, while allowing $k>1$ to average over multiple plausible masked-context views of $\mathbf{y}$}.
$f_{\theta}$’s effect on PRISM. We show in Proposition 1 and Proposition 2 that, as intended, the design of (b), in which we use $f_{\theta}$ to obtain $\mathbf{y}$}, does not directly influence our theoretical results. To clarify, when $k=1$, the minimizer $g_{\phi^{\star}}^{i}({\mathbf{y}})\;=\;p\!\left({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,\middle|\,{\mathbf{y}}\oplus\mathbf{m}_{i}\right)$ remains fully independent from $f_{\theta}$ (or the rule for choosing $i$), and when $k>1$, the only change is indirect: $f_{\theta}$ (and the rule for choosing index set $\mathcal{S}$ during PRISM fine-tuning) affects the posterior over masked sets $\Pi(\cdot\mid{\mathbf{y}},i)$.
@@ Appendix C Employing PRISM at inference
9:  Sample $\mathcal{S}\subset\mathcal{M}$ according to the unmasking rule
10:  Sample $\mathcal{T}\subset\{i\mid{\mathbf{x}}_{t_{\ell}}^{i}\neq\mathbf{m}\}$ with low-$|\mathcal{T}|$ indices of $\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{\cdot}(x_{t_{\ell}})$}
11:  Unmask ${\mathbf{x}}^{i}_{t_{\ell}}$ to $v^{i}\sim{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{x}}_{t_{\ell}})}$ for each $i\in\mathcal{S}$
12:  Remask ${\mathbf{x}}^{j}_{t_{\ell}}$ for each $j\in\mathcal{T}$
13:  ${\mathbf{x}}_{t_{\ell+1}}\leftarrow{\mathbf{x}}_{t_{\ell}}$
14:return ${\mathbf{x}}_{t_{N}}={\mathbf{x}}_{0}$
@@
9:  Sample $\mathcal{S}\subset\mathcal{M}$ according to the unmasking rule
10:  Sample $\mathcal{T}\subset\{i\mid{\mathbf{x}}_{t_{\ell}}^{i}\neq\mathbf{m}\}$ with low-$|\mathcal{T}|$ indices of $\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{\cdot}(x_{t_{\ell}})$}
11:  Unmask ${\mathbf{x}}^{i}_{t_{\ell}}$ to $v^{i}\sim{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{x}}_{t_{\ell}})}$ for each $i\in\mathcal{S}$
12:  Remask ${\mathbf{x}}^{j}_{t_{\ell}}$ for each $j\in\mathcal{T}$
13:  ${\mathbf{x}}_{t_{\ell+1}}\leftarrow{\mathbf{x}}_{t_{\ell}}$
14:return ${\mathbf{x}}_{t_{N}}={\mathbf{x}}_{0}$
C.1 PRISM-Loop
As we mentioned in Section 4.2, the sentence diversity is crucial for unconditional generation, we observe that the vanilla PRISM inference from Section 3.3 can degrade diversity. we therefore employ PRISM-loop, analogous to ReMDM-loop (wang2025remasking), where we extend the standard generation process with a terminal refinement phase. After the initial generation of a sequence, an additional $N_{l}$ looping steps are performed. In each of these steps, a fixed-size subset of $M$ tokens is randomly selected from the full sequence, remasked, and subsequently unmasked to allow for iterative refinement.
Algorithm 4 PRISM inference with loop strategy
1:Require: Fine-tuned MDM with unmasking posterior model $f_{\theta}$, per-token quality head $g_{\theta}$, sampling steps NN, sequence length $L$, discretized time steps $\{t_{\ell}\}_{\ell=0}^{N}$, remasking token ratio $\eta>0$, loop steps $N_{l}$, number of remasking tokens at the loop stage $M$
2:Sample ${\mathbf{x}}_{0}$ from ${\mathbf{x}}_{1}$ for NN steps with PRISM with $\eta$
3:for $l=1$ to $N_{l}$
4:  Uniformly sample size $M$ subset $\mathcal{T}\subset\{0,\dots,L-1\}$
5:  Remask ${\mathbf{x}}^{i}_{0}$ for each $i\in\mathcal{T}$
6:  Unmask ${\mathbf{x}}^{i}_{0}$ to $v^{i}\sim{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{x}}_{0})}$ for each $i\in\mathcal{T}$
7:return ${\mathbf{x}}_{0}$

### v4

@@ 4.3 Applying PRISM to 8B MDM
Setup. We take LLaDA-88B-Instruct (Nie et al., 2025), an open-sourced 88B MDM. While freezing the backbone, we add an additional head and a LoRA adapter (Hu et al., 2022), for a total of $\approx 250$M trainable parameters. We fine-tune it using the PRISM loss on the opc-sft-stage2 (Huang et al., 2024) (0.10.1M pairs) for $100$ epochs, which takes $\approx 30$ hours on $12$ NVIDIA H100 GPUs.
@@
Results. We evaluate the resulting model in a zero-shot setting on MBPP (Austin et al., 2021b) and HumanEval (Chen, 2021), a challenging Python coding benchmarks. As shown in Table 1, PRISM outperforms ReMDM and ReMDM-conf across different numbers of sampling steps. We detail training and inference configurations in Appendix D.8D.9.
4.4 Mechanistic Analysis of PRISM
In this section, we investigate PRISM’s behavior at a fine-grained level, showing that the overall PRISM pipeline operates as desired.
Calibration analysis. We measure how well the learned per-token quality scores are calibrated to the targeted likelihood. Using the fine-tuned model from Section 4.2, we collect per-token quality scores $g_{\theta}^{i}(\mathbf{x}_{t})$ for all inference-time sequences $\mathbf{x}_{t}$s and their clean positions $i$. Then we also compute the corresponding targeted likelihood queried from $f_{\theta}$, namely $f_{\theta}^{i}(\mathbf{x}_{t}^{i}\mid\mathbf{x}_{t}\oplus\mathbf{m}_{i})$. Surprisingly, the resulting calibration curves (Figure 3.(b); additional results deferred to Figure 6 in the appendix) closely track the queried likelihoods, indicating that $g_{\theta}^{i}(\mathbf{x}_{t})$ closely calibrates the target posterior.
Error-correction analysis. To assess the actual error-correction behavior of PRISM, we take the model in Section 4.3 and measure how often it corrects the syntax errors. Across 164 problems in HumanEval, PRISM corrects 19 out of 26 syntax errors. This provides fine-grained evidence that PRISM can identify and remediate errors in practice. For clarity, we report several failure and success instances in Appendix D.7.
PRISM-tuned model’s knowledge transfers to another. As we underscore in Section 3.1, the ground-truth per-token quality is defined with respect to the data distribution and thus does not depend on a particular pretrained model. This suggests that a PRISM fine-tuned MDM $A$ can be applied to another pretrained MDM $B$ as a self-corrector, given they operate over the same distribution, e.g., Sudoku puzzles, coding.
As a proof-of-concept experiment on this idea, we apply our PRISM fine-tuned model in Section 4.3 as a self-corrector while using LLaDA-8B-Instruct (not fine-tuned) as a backbone unmasking model. On HumanEval under the same evaluation protocol, self-correction yields a substantial gain of $10.9\%$ (from $23.2\%$ to $34.1\%$). This indicates that the PRISM-tuned model can serve as a general-purpose self-corrector for other pretrained MDMs.
Learning likelihood itself vs. scalar. PRISM learns a scalar per-token quality score, rather than the full per-position likelihood $p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i})$. The main motivation is training efficiency: as a scalar score is already informative for self-correction, and learning a full categorical distribution can be computationally inefficient.
The PRISM framework can also be extended to likelihood learning by replacing the BCE objective in Equation 2 with a cross-entropy objective. This richer supervision can distinguish genuinely incorrect tokens from high-entropy positions with many plausible alternatives, enabling uncertainty-aware remasking at additional training cost. We provide additional discussion in Appendix D.6.
@@ 5 Conclusion
Discussion on experimental results. MDMs model per-position unmasking posteriors rather than the joint distribution over positions. Consequently, with few-step inference, they are prone to dependency errors across positions (Liu et al., 2024a; Xu et al., 2024; Kang et al., 2025), which creates a natural opportunity for our self-correction mechanism to detect erroneous predictions. Indeed, the performance gap between PRISM and the vanilla inference increases as the number of sampling steps decreases (Figure 3, Table 1).
@@
A unique feature of the PRISM loss is that it provably learns per-token quality without fully generating a clean sequence. This sheds light on PRISM’s potential compared to other approaches, such as reinforcement learning, where full sequence generation is necessary to obtain a learning signal.
That said, our notion of per-token quality score has limits; as it is based on the per-position posterior marginals, it cannot fully capture global errors, e.g., reasoning. To our knowledge, nevertheless, no prior work principally even targets the per-token quality within the efficient training objective. We therefore view this work as a step that opens the door to developing frameworks that enable discrete diffusion models to correct global reasoning errors.
@@
Impact Statement
This paper advances understanding of discrete diffusion models and contributes to the broader machine learning literature. We do not anticipate societal impacts that warrant specific discussion beyond these general considerations.
@@ Appendix A Related Works
A.1 Background and Related Literature
Masked Diffusion Models. Masked Diffusion Models (MDMs) were previously built on continuous-time Markov chains over discrete spaces (Hoogeboom et al., 2021). Subsequently, Austin et al. (2021a) introduced D3PM with several families of transition kernels, followed by SEDD (Lou et al., 2023). These discrete diffusions can also be understood through the lens of discrete flows (Campbell et al., 2024; Gat et al., 2024), which parallel flow matching and stochastic interpolants over continuous spaces (Albergo and Vanden-Eijnden, 2022; Albergo et al., 2023; Lipman et al., 2022). Among transition-kernel designs, the absorbing state (masking) kernel has been particularly popular due to its simple formulation and training objective. Subsequent work established a theoretical framework for these models (Zheng et al., 2023; Sahoo et al., 2024; Zheng et al., 2024). Follow-up works have scaled MDMs up to billions of parameters, showcasing their potential compared to autoregressive models (Nie et al., 2025; Ye et al., 2025; DeepMind, 2025; Labs et al., 2025; Gong et al., 2025).
Inference-time strategies for discrete diffusion models. A key feature of MDMs is their inference-time flexibility, highlighted by Zheng et al. (2024); Ou et al. (2024) and aligned with our any-order perspective in Section 2.1. Recall the MDM inference procedure from Section 2.1:
@@ A.1 Background and Related Literature
The choice of $\mathcal{S}$ is entirely flexible, and an appropriate selection can substantially improve downstream task performance. Since our work focuses on self-correction, i.e., re-masking, we only briefly survey prior work on unmasking strategies (Kim et al., 2025; Peng et al., 2025; Chang et al., 2022; Ben-Hamu et al., 2025; Rout et al., 2025). A complementary line of work studies inference-time planning strategies for discrete diffusion with uniform transition kernels (Liu et al., 2024b; Varma et al., 2024); these methods fall outside the scope of MDMs.
A.2 Details on prior work on remasking strategy (self-correction) for MDM
@@
  • DFM (Gat et al., 2024) and ReMDM (Wang et al., 2025) randomly remask clean tokens in $\mathbf{y}$}, which is equivalent to assigning i.i.d. per-token qualities $g_{\star}^{i}({\mathbf{y}})\sim\operatorname{Unif}[0,1]$.
  • P2-Self (Peng et al., 2025) defines per-token quality using the pretrained MDM’s score for the observed token under input $\mathbf{y}$}, e.g., the probability $f_{\theta}^{i}({\mathbf{y}}^{i}\mid{\mathbf{y}})$. Given that ${\mathbf{y}}^{i}\neq\mathbf{m}$, the model was never trained on this index in this state; it only encodes a meaningful quantity (unmasking posterior) when ${\mathbf{y}}^{i}=\mathbf{m}$. This notion of per-token quality is therefore not grounded.
  • ReMDM-conf (Wang et al., 2025) takes the per-token quality to be the likelihood at the step the token was unmasked: if position $i$ was unmasked at $t_{k}>t_{\ell}$ (i.e., an earlier step), set $g_{\star}^{i}:=f_{\theta}^{i}({\mathbf{x}}_{t_{k}}^{i}\mid{\mathbf{x}}_{t_{k}})=f_{\theta}^{i}({\mathbf{x}}_{t_{\ell}}^{i}\mid{\mathbf{x}}_{t_{k}})$. As noted, this ignores the current state ${\mathbf{x}}_{t_{\ell}}$ and can therefore be misaligned with the model’s present beliefs.
  • Token Critic (Lezama et al., 2022a, b) and P2-train (Peng et al., 2025) train an auxiliary scorer for the likelihood of a given prediction of clean sequence $\hat{{\mathbf{x}}}$. Given ${\mathbf{x}}_{t_{\ell}}$, they first form a full completion $\hat{{\mathbf{x}}}$ by unmasking each mask token at position $i$ from $f_{\theta}^{i}(\cdot\,\mid\,{\mathbf{x}}_{t_{\ell}})$, then evaluate quality by passing this $\hat{{\mathbf{x}}}$ to the external module. A key limitation is that $\hat{{\mathbf{x}}}$ is drawn from per-token unmasking posteriors, not the ground-truth posterior across positions; thus, even with a perfect pretrained MDM, $\hat{{\mathbf{s}}}$ can deviate substantially from the true likelihood ${\mathbf{x}}\sim p_{\mathrm{data}}$, losing a theoretical guarantee.
  • GIDD (von Rütte et al., 2025) introduces a family of transition kernels that linearly combine uniform and absorbing (masking) kernels. Models pretrained under this kernel encode posteriors at clean-token positions–however, this requires training under the new kernel and is not directly applicable to an off-the-shelf pretrained MDM, altogether losing the benefit of MDM over other types of kernels.
  • Informed corrector (Zhao et al., 2024) employs a hollow-transformer $g_{\phi^{\prime}}$ satisfying $g_{\phi^{\prime}}^{i}({\mathbf{y}})=g_{\phi^{\prime}}^{i}({\mathbf{y}}\oplus\mathbf{m}_{i})\approx p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\mid{\mathbf{y}}\oplus\mathbf{m}_{i})$, enabling the same training loss as MDM. This, however, mandates a modified architecture (attention tweak) that processes $\mathbf{y}$} as if position $i$ were masked, necessitating retraining.
@@ B.2 Marginal Equivalence of the Masking Process
We formalize that the two masking procedures stated in Section 2.1 induce the same conditional distribution of $\mathbf{z}$} given $\mathbf{x}$}, hence the same joint distributions $({\mathbf{x}},{\mathbf{z}})$. It suffices to show they generate the same distribution over masked index sets. Consider (b): draw $t\sim\mathrm{Unif}[0,1]$ and for each $i$, replace ${\mathbf{x}}^{i}\in\mathcal{V}$ with $\mathbf{m}$ with probability $t$. The probability of $n$ tokens to be masked is:
@@
where $\mathrm{B}$ denotes the Beta function. This coincides with procedure (a), where $n$ is uniformly sampled from $\{0,1,\dots,L\}$.
@@ B.3 PRISM provably learns the per-token quality
In this section, we restate Proposition 3.1 and detail its proof. The proof essentially follows Section 3.1, but we provide the detailed calculations here.
Proposition 3.1 (restated). Let the PRISM loss be
@@ B.4 Extension on PRISM’s provable guarantee
In this section, we provide the pseudocode of the practically used PRISM fine-tuning pipeline and then state its theoretical guarantee. We note that the choice of index set $\mathcal{S}$ at line 5 of Algorithm 1 is flexible; it can either be random or confidence-based. To clarify, at line 6, although we use $f_{\theta}$ to obtain $\mathbf{y}$}, we use the stop-gradient operation.
@@
Proposition B.1 (Provable guarantee of PRISM-extension).
@@
By using this property, we now complete the proof; the remaining parts are the same as Proposition 3.1.
@@
Discussion on a new notion of per-token quality. Proposition B.1 guarantees that the minimizer of the extended PRISM loss satisfies
@@
Note that by construction, we always have $i\in\mathcal{S}$. This encompasses the $k=1$ case in Section 3.1; When $k=1$, the posterior distribution $\Pi(\cdot\,|\,{\mathbf{y}},i)$ collapses to a point mass at $\mathcal{S}=\{i\}$, recovering Proposition 3.1.
We emphasize that this extended definition still captures our intended notion of per-token quality—namely, the (posterior) likelihood of ${\mathbf{y}}^{i}$ given the rest of the context, here defined under a certain distribution over ${\mathbf{y}}\oplus\mathbf{m}_{\mathcal{S}}$. Consequently, a model fine-tuned with the (extended) PRISM loss retains the capacity to identify low-quality tokens, while allowing $k>1$ to average over multiple plausible masked-context views of $\mathbf{y}$}.
$f_{\theta}$’s effect on PRISM. We show in Proposition 3.1 and Proposition B.1 that, the design of (b), in which we use $f_{\theta}$ to obtain $\mathbf{y}$}, does not directly influence our theoretical results. To clarify, when $k=1$, the minimizer $g_{\phi^{\star}}^{i}({\mathbf{y}})\;=\;p\!\left({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,\middle|\,{\mathbf{y}}\oplus\mathbf{m}_{i}\right)$ remains fully independent from $f_{\theta}$ (or the rule for choosing $i$), and when $k>1$, the only change is indirect: $f_{\theta}$ (and the rule for choosing index set $\mathcal{S}$ during PRISM fine-tuning) affects the posterior over masked sets $\Pi(\cdot\mid{\mathbf{y}},i)$.
@@ Appendix C Employing PRISM at inference
9:  end if
10:  Sample $\mathcal{S}\subset\mathcal{M}$ according to the unmasking rule
11:  Sample $\mathcal{T}\subset\{i\mid{\mathbf{x}}_{t_{\ell}}^{i}\neq\mathbf{m}\}$ with low-$|\mathcal{T}|$ indices of $\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{\cdot}(x_{t_{\ell}})$}
12:  Unmask ${\mathbf{x}}^{i}_{t_{\ell}}$ to $v^{i}\sim{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{x}}_{t_{\ell}})}$ for each $i\in\mathcal{S}$
13:  Remask ${\mathbf{x}}^{j}_{t_{\ell}}$ for each $j\in\mathcal{T}$
14:  ${\mathbf{x}}_{t_{\ell+1}}\leftarrow{\mathbf{x}}_{t_{\ell}}$
15:end for
16:return ${\mathbf{x}}_{t_{N}}={\mathbf{x}}_{0}$
@@
9:  end if
10:  Sample $\mathcal{S}\subset\mathcal{M}$ according to the unmasking rule
11:  Sample $\mathcal{T}\subset\{i\mid{\mathbf{x}}_{t_{\ell}}^{i}\neq\mathbf{m}\}$ with low-$|\mathcal{T}|$ indices of $\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{\cdot}(x_{t_{\ell}})$}
12:  Unmask ${\mathbf{x}}^{i}_{t_{\ell}}$ to $v^{i}\sim{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{x}}_{t_{\ell}})}$ for each $i\in\mathcal{S}$
13:  Remask ${\mathbf{x}}^{j}_{t_{\ell}}$ for each $j\in\mathcal{T}$
14:  ${\mathbf{x}}_{t_{\ell+1}}\leftarrow{\mathbf{x}}_{t_{\ell}}$
15:end for
16:return ${\mathbf{x}}_{t_{N}}={\mathbf{x}}_{0}$